# GPS-Synchronized Monitoring of Supernova Bursts in PandaX-4T: Enabling a Neutrino Trigger for Multi-Messenger Astronomy Binyu Pang

### v1

@@
GPS-Synchronized Monitoring of Supernova Bursts in PandaX-4T: Enabling a Neutrino Trigger for Multi-Messenger Astronomy
@@
The landmark detection of neutrinos from SN 1987A marked the dawn of a new era in astrophysics, unequivocally establishing them as indispensable probes for investigating stellar evolution and supernova dynamics. Today, supernova burst neutrinos can be observed through coherent elastic neutrino-nucleus scattering in tonne-scale liquid xenon detectors designed for direct dark matter detection. Capitalizing on this capability, we have developed and deployed a real-time supernova monitoring system for the PandaX-4T experiment. This system incorporates a GPS module providing millisecond-level timing precision, achieves a low false-alarm rate, and maintains high sensitivity to galactic supernova events. This robust methodology is directly scalable and slated for implementation in the next-generation PandaX-20T experiment.
Core-collapse supernova explosions occur fewer than three times per century within the Milky Way [1]. The most recent neutrinos from a supernova (SN) explosion were detected from SN 1987A, with around 20 neutrino events in total observed by the global neutrino detectors [2, 3, 4]. The mechanism of supernova explosions has been explained by a series of theoretical models [5, 6, 7]. However, given the limited statistics, scientists have not yet fully understood the physical processes in supernova dynamics yet. To get prepared for the neutrino detection from next galactic supernova burst, a number of neutrino experiments have implemented specialized supernova monitoring systems [8, 9, 10].
@@
Supernova bursts occur in the latter stages of evolution of massive stars with masses exceeding eight times that of the Sun. Roughly erg of energy is released when a SN bursts, with neutrinos carrying away about of the energy [11]. The entire explosion process takes about several seconds. Compared to electromagnetic radiations and gravitational waves, neutrinos generated in the stellar core can more easily escape the supernova interior and are emitted earlier, enabling them to reach ground-based observatories before the light and gravitational messengers [12]. Therefore, a global organization known as the Supernova Early Warning System (SNEWS) was established to provide the astronomical community with an early warning of SN explosions [13, 14].
The COHERENT experiment reported the first observation of coherent elastic neutrino nucleus scattering (CENS) using neutrinos from a spallation neutron source in 2017 [15]. This observation provides an excellent opportunity for detecting astronomical neutrinos by CENS [16, 17]. The CENS cross-section is proportional to the square of the target nuclei’s neutron number, , where is the neutron number of the target nucleus, is the Fermi coupling constant, and denotes the neutrino energy [18]. Because the xenon target nucleus contains around 80 neutrons, it is a suitable choice for a direct detection experiment of dark matter utilizing liquid xenon, such as XENONnT/LUX-ZEPLIN [9, 19], to detect SN neutrinos by CENS. The neutral current process makes this interaction sensitive to all neutrino flavors, and the neutrino oscillation is not necessary to be taken into account. It can also complement the findings from other experiments that are particularly sensitive to only one or several specific flavors of neutrinos [20, 21]. On the other hand, detecting neutrinos via CENS is seriously hampered by the low nuclear recoil energy of keV level.
@@
The PandaX-4T experiment is located in the China Jinping Underground Laboratory (CJPL), with 2,400 m of rock overhead. Comprehensive knowledge about the detector can be found in Ref. [22], and only a few essential details are described in the Letter. The primary detector is a cylindrical xenon time projection chamber (TPC), containing approximately 3.7 tonnes of liquid xenon (LXe). Physical events are reconstructed according to the coincidence detection of photons from prompt scintillation (S1) and delayed electroluminescence photons from ionized electrons (S2). The top PMT pattern and the time difference between S1 and S2 are used to reconstruct the 3D interaction vertex, which can further suppress the external background through fiducialization.
Precise measurement of signal arrival times in the detector is crucial for tracking the SN by time-delay triangulation, which requires combining data from several observatories. A millisecond-level GPS-based timing system was specially designed for PandaX-4T, as shown in Fig. 1, to facilitate collaborative early warnings of SN explosions with other neutrino experiments globally. The GPS ground sub-system comprises a GPS receiver and a camp control module. The GPS antenna receives satellite signals and transmits them to the GPS server, which generates a high-precision 10 MHz reference clock through its internal clocking mechanism to synchronize the local clock of the control module. The precise GPS time, encoded in the IRIG-B code and synchronized with the one pulse per second (PPS) signal, is transmitted to the control module in the underground laboratory (hereafter referred to as lab control module) through a multi-kilometer optical fiber network. A hardware clock counter of the lab control module synchronizes with the DAQ’s master clock at the start of each new data acquisition run, and the mapping between the delivered GPS time and the hardware clock ticks is stored in a database. The time delay introduced by the fiber optic length is determined by measuring the time difference between the emission and reception of synchronized time signals (e.g., 1 PPS pulses) between the camp control module and the lab control module. This delay is utilized to correct timestamps during data acquisition. The precision of this system is maintained at the millisecond level, which is determined primarily by the GPS server.
@@
The PandaX-4T has been in operation since November 2020 and is scheduled to continue the data taking until the end of 2025. To improve the detection time accuracy of SN explosions, a dedicated GPS-based timing system as described above was established in July 2025.
Figure 1: Schematic diagram of the PandaX-4T’s GPS-based timing system.Figure 2: Schematic diagram of SN online monitoring system in PandaX-4T.
Using the transient increase in detector signals induced by neutrinos from supernova explosions, PandaX-4T has developed an online software-based monitoring system to identify SN events, as illustrated in Fig. 2. When the DAQ system writes a new raw data file, the acquired data is subsequently stored utilizing the Bamboo Shoot3 library [23]. This library offers capabilities for custom formatting and compression of commonly used data structures, enabling efficient data management and storage [24]. Each raw file contains roughly one minute of data and is approximately 2 GB in size. The data transformation process takes about ten seconds and compresses the data size to approximately 1 GB. Then, the data is transmitted to the remote data center in Chengdu for offline processing and reconstruction, which takes about 5 minutes. After event reconstruction, information of each recorded event, including interaction vertex coordinates, S1 and S2-related variables, etc., is derived. Next, the SN neutrino cut, including the dead time cut which refers to a period of time removed after large signals, the fiducial volume (FV) cut, the accidental coincidence (AC) cut, and quality cuts for S1 and S2, etc., is applied to further reduce the background and the surviving events serve as SN neutrino candidates to investigate false alert rates. The monitoring results indicate that the background event rate is stable. Finally, the SN trigger algorithm is utilized to identify potential SN explosion events. An alert will be generated when the number of neutrino candidates exceeds the threshold () within a 10-second time window. Further details on this method can be found in Ref. [25]. The SN warning information, including the start time of the alert and the total number of neutrino candidates, etc., will be sent to the PandaX-4T SN group via email. It takes approximately 5 seconds to complete the process from selecting SN neutrino candidates to sending out an alert. The operational status of the SN monitoring system can be displayed on a webpage, allowing relevant personnel to inspect key parameters (e.g., event rate). Accounting for the aforementioned latency factors, the entire monitoring process takes approximately 6 minutes. The system sustains stable operation during routine data-taking periods, except in exceptional circumstances, such as power outages or server maintenance. In addition, during the calibration process, the SN monitoring program automatically excludes calibration-related data based on its data type.
The background in the detector is almost negligible over a time window of a few seconds. Practically, the background event rate in PandaX-4T is estimated to be approximately after the supernova neutrino cuts, resulting in a false alert rate of about twice per month. The majority of these backgrounds are electron recoil (ER) events, which, similar to previous analyses [26], comprise the material radioactivity, environmental radioactivity, AC events, and xenon contaminants (e.g., 222Rn). In about 83 days of physics data, six alerts are expected and eight are observed, which are in good agreement. Further verification of the consistency between the expectation and the observation is carried out with neutron calibration data from deuteron-deuteron (DD) and AmBe. The results show that 886 (288) alerts are observed for approximately 2.8 (3.7) days of the DD (AmBe) data, compared to 870.6 (265) expected alerts, corresponding to the background event rate of roughly () . The algorithm of the expectation can be found in Ref. [25]. Regarding DD and AmBe data, scenarios where calibration event candidates are present in two separate files are taken into account, while for physics data, this can be ignored due to the lower background event rate. In the future, the false alert rate can be further reduced by lowering the background level, including using low-radioactivity materials in the detector and enhancing the removal efficiency of radioactive elements through a distillation tower.
@@
Detection efficiency can be obtained via the pseudo data based on the simulation and assembled” real detector waveforms and verified through calibration data, which is the same as the approach adopted in Refs. [27, 28, 26]. The signal model for PandaX-4T is built with the Noble Element Simulation Technique (NEST) v2.3.6 parameterization [29], and the relevant parameters are tuned by fitting calibration data [30]. The NEST model can simulate the light yield (, number of photons per keV) and charge yield (, number of electrons per keV) of nuclear recoil events in LXe. Waveform simulation (WS) is developed using a data-driven simulation program in PandaX-4T, which generates complete waveforms of signals and background using and as inputs [31]. The simulated events are reconstructed using the same methods as the experimental data.
The overall efficiency comprises the efficiencies of signal reconstruction, data quality selection, and region-of-interest (ROI). The signal reconstruction efficiency is similar to that discussed in the previous analysis [25]. Data quality selection mentioned earlier is inherited from the previous WIMP analysis [26], with slight adjustments made to adapt to the current data set. The ranges of S1 and S2 charges in the ROI are changed to [2.1, 100] PE and [80, 4000] PE to balance the false alert rate due to background contributions and detection efficiency of the nuclear recoil events induced by SN neutrinos. The efficiency of the dead time cut is estimated to be , representing the surviving probability of the live time. The overall detection efficiency of SN neutrino events using the Garching model of described in the following paragraph is .
Figure 3: Typical timing distributions of the detected supernova neutrino events from the distance of 168 pc (top left) and 10 kpc (top right). Data points are produced by simulations and the red lines are the fit results. The fitted onset times relative to the truth time in 10,000 simulated measurements are shown in the bottom panels. The resulting onset time measurement exhibits a variation of ms and s relative to zero for 168 pc and 10 kpc, respectively, within a 68% confidence level (C.L.).
The electroweak neutral current process allows neutrinos to coherently interact with all nucleons in the nucleus when their energy is very low ( MeV). The differential cross-section calculation is detailed in Ref. [25]. To investigate the nuclear recoil signals induced by supernova neutrinos in PandaX-4T, supernova models are used to obtain the neutrino flux as input. In this work, we adopt two Garching models with and , employing the LS220 nuclear equation of state (EoS) [32, 33]. More information is available in the earlier publication [25]. Subsequently, the differential event rate of CENS can be expressed as
where represents the time after the SN core bounce, tonnes is the effective target mass in this analysis, is Avogadro’s number, and the molar mass of xenon is , and is the distance from the SN to Earth. represents the minimum neutrino incident energy required to induce a nuclear recoil energy of in CENS. In practice, we integrate over MeV). and are the differential cross-section of the CENS process and the total neutrino fluxes of all flavors at the energy of and time of , respectively. represents the detection efficiency.
The core bounce onset time can be determined when sufficient event statistics are available. A typical fitting result of the timings after the core bounce of the supernova neutrino events from the distance of 168 pc is displayed in the top left panel of Fig. 3. The points represent the simulated SN neutrino events, and the histogram represents the fitting result using the equation , where f denotes the differential time profile of the SN events, while a and b signify the amplitude and time shift of the time profile, respectively, which are free parameters. The bottom left panel of Fig. 3 illustrates the distribution of parameter b, indicating a discrepancy in the core bounce time of SN explosions between the simulated measurements and the theoretical predictions. The case of 10 kpc is shown in the right panels of Fig. 3. The 68% C.L. intervals for the variation of the offset times relative to the zero are ms at 168 pc and s at 10 kpc due to smaller statistics.
The expected number of SN neutrino events in the detector can be calculated by integrating Eq. 1. Here, we integrate the variable for the model with ( and ) over the time interval from the onset time of core bounce to 8.35 (7.6) seconds. is integrated up to 30 keV. The results indicate that the number of the expected SN neutrinos from 10 kpc (168 pc) in PandaX-4T is approximately 8.1 () and 3.7 () for the model of and , respectively. Only the results for the case of is shown in Fig. 3 for simplicity. The distances of 10 kpc and 168 pc correspond to the positions of the nearby Galactic center and Betelgeuse [34] as observed from Earth, respectively.
Figure 4: Detection probability of the SN explosions as a function of distance using the Garching model with for PandaX-4T (black) and PandaX-20T (red). The corresponding false alert rate is twice a month in PandaX-4T.
The detection probability of SN explosions is estimated using the Garching model with , employing the LS220 EoS. Figure 4 illustrates the detection probability as a function of SN distance to the PandaX-4T and PandaX-20T. PandaX-4T achieves nearly- detection efficiency at 10 kpc, while the PandaX-20T with the effective target mass enlarged by a factor of about 6 to tonnes, maintains 100% efficiency up to 30 kpc.
SNEWS significantly improves the confidence of SN explosion warnings through coincident onset times of alerts from several collaborative experiments. The SNEWS system provides different message tiers for participating experiments, as listed in the first column of Table 1. The second and third columns present the essential information for each tier and the corresponding values of a typical SN candidate alert in the PandaX-4T, respectively. The coincidence tier is a prerequisite for experiments to be incorporated in SNEWS. More details about SNEWS alert software are provided in Ref. [35]. The detector status parameter is set to ON/OFF according to its operational condition. The neutrino time is represented in UTC format, derived from GPS time in Pandax-4T. The time series is an ordered sequence of nanosecond-precision integers, representing the time offsets relative to the initial neutrino time. According to the background-only assumption, if two events are observed in the first 5-second time window, the p-value calculated via the Poisson distribution with expected rate (where and s) is . In contrast, the p-value for observing greater than or equal to zero events is 1 in the second 5-second time window. Retraction is initiated via a dedicated command provided by the SNEWS network, based on real-time experimental considerations. The information shown in the third column of Table 1 satisfies the requirements of the SNEWS message tiers.
Table 1: List of different message tiers within the SNEWS network. The detector name is required in each tier and is not listed in the column. The neutrino time is represented in UTC format. The p-value is calculated via the Poisson distribution for every 5-second time window. The units for the time bins and the neutrino time series (measured relative to the first event) are seconds and nanoseconds, respectively.
@@
In summary, we have established an online supernova monitoring system with GPS time implemented at PandaX-4T. The monitoring system can provide the astronomical community with early warning of SN explosions and can achieve about detection efficiency of SN explosions at 10 kpc in PandaX-4T, corresponding to a false alert rate of twice per month. Distinguishing between different SN models would be possible using the distinct time profile and energy spectrum features if an SN burst were to occur at a close distance to Earth (e.g., Betelgeuse). In the near future, the PandaX-20T experiment is expected to significantly enhance the sensitivity to astronomical neutrinos, particularly those from core-collapse supernova in our galaxy and the Sun, thereby enabling more precise tests of astrophysical models.
@@
Acknowledgements. This project is supported in part by grants from National Key R&D Program of China (Nos. 2023YFA1606200, 2023YFA1606203), National Science Foundation of China (Nos. 12090060, 12090061, U23B2070), and by Office of Science and Technology, Shanghai Municipal Government (grant Nos. 21TQ1400218, 22JC1410100, 23JC1410200, ZJ2023-ZD-003). We thank for the support by the Discipline Construction Fund of Shandong University, and the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Thomas and Linda Lau Family Foundation, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.

### v2

# GPS-Synchronized Monitoring of Core-collapse Supernova Bursts with PandaX-4T via Coherent Elastic Neutrino Nuclear Scattering Binyu Pang

@@
GPS-Synchronized Monitoring of Core-collapse Supernova Bursts with PandaX-4T via Coherent Elastic Neutrino Nuclear Scattering
@@
The landmark detection of neutrinos from SN1987A marked the dawn of neutrino astrophysics. The neutrino burst provided essential insights into fundamental properties of neutrinos, and served as key probes of stellar evolution and supernova dynamics. The recent advancement in coherent elastic neutrino-nucleus scattering enables the detection of core-collapse supernova burst neutrinos using tonne-scale liquid xenon detectors originally designed for dark matter direct detection. Leveraging this capability, we developed and deployed an online supernova monitoring system for the PandaX-4T experiment. This system features a GPS module with millisecond-level timing precision, a low false-alarm rate, and high sensitivity to galactic core-collapse supernova explosion events. The methodology is robust, directly scalable, and planned for implementation in the next-generation PandaX-20T experiment.
@@
The core-collapse supernova (CCSN) explosions occur fewer than three times per century within the Milky Way [1]. The CCSN bursts occur in the latter stages of evolution of massive stars with masses exceeding eight times that of the Sun. Roughly erg of energy is released during a CCSN burst, with neutrinos carrying away about of the energy [2]. The neutrino burst lasts only a few or a few tens of seconds, and unlike electromagnetic radiation or gravitational waves, neutrinos can escape the stellar core more easily and are emitted earlier, allowing them to reach ground-based observatories before other messengers [3]. The only supernova (SN) neutrinos ever detected were from SN1987A, with around 20 events observed by a few detectors [4, 5, 6]. These neutrino events provided critical insights into the fundamental properties of neutrinos, and inspired theoretical models on CCSN explosion mechanisms [7, 8, 9]. However, progress in the field remains hindered by the lack of statistics of SN neutrinos. To prepare for neutrino detection from the next galactic CCSN burst, several neutrino experiments have implemented specialized supernova monitoring systems [10, 11, 12, 13], forming a global organization known as the Supernova Early Warning System (SNEWS) [14, 15].
Recently, significant advancements have been made in neutrino detection methods. In 2017, with neutrinos from a spallation neutron source, the COHERENT experiment reported the first observation of coherent elastic neutrino-nucleus scattering (CENS) [16], a coherently enhanced weak interaction predicted by Freedman in 1974 [17]. CENS offers a remarkable opportunity for detecting low energy astrophysical neutrinos [18, 19, 20, 21, 22], as its cross-section scales approximately with the square of the target nucleus’s neutron number, , where is the neutron number, is the Fermi coupling constant, and is the neutrino energy [23]. Furthermore, as a neutral current process, CENS is sensitive to all neutrino flavors, providing a valuable complement to experiments that primarily focus on specific flavors [24, 25]. The challenge, however, lies in the extremely low nuclear recoil energy, which is difficult to detect. Over the past two decades, liquid xenon-based dark matter direct detection experiments, designed to detect rare and low-energy scattering events, have made remarkable progress in improving their sensitivities. Experiments such as XENONnT, LUX-ZEPLIN (LZ), and XMASS [11, 26, 27] have proposed leveraging CENS to detect SN neutrinos. Notably, in 2024, the PandaX-4T [21] and XENONnT [22] collaborations independently reported indications of CENS from solar 8B neutrinos with xenon nuclei, and very recently, the LZ collaboration also reported the evidence of 8B CENS with more than 4- significance [28]. These findings mark a significant milestone towards the realization of SN neutrino detection via CENS.
In this paper, we will discuss the performance of a newly developed CCSN burst monitoring system in PandaX-4T,
utilizing online selection of CENS events.
@@
The PandaX-4T experiment is located in the China Jinping Underground Laboratory (CJPL), shielded by 2,400 meters of rock overburden. Comprehensive details about the detector can be found in Ref. [29], with only a few essential aspects highlighted in this Letter. The primary detector is a cylindrical xenon time projection chamber (TPC) containing approximately 3.7 tonnes of liquid xenon (LXe). A physical event produces a delayed coincidence of photons from prompt scintillation (S1) and delayed electroluminescence photons from ionized electrons (S2), which are detected via the top and bottom photomultiplier (PMT) arrays. The waveforms from all fired PMTs are read out under a continuous and triggerless” data acquisition mode, and event building and reconstructions are performed by analysis software. The three-dimension interaction vertex is reconstructed using the top PMT pattern and the time difference between S1 and S2, allowing for further suppression of external background through fiducialization.
@@
The PandaX-4T has been in operation since November 2020 and is scheduled to continue the data taking until the end of 2025. To allow an online monitoring for potential CCSN explosions, a dedicated GPS-based CCSN burst system was established in July 2025. About 83 days of physics data collected between July 18 and October 14, 2025, will be used to evaluate its performance.
A millisecond-level GPS-based timing system was specially designed for PandaX-4T, as shown in Fig. 1, to facilitate collaborative early warnings with other neutrino experiments via time-delay triangulation [15, 30]. The GPS ground sub-system comprises a GPS receiver and a camp control module. The GPS antenna receives satellite signals and transmits them to the GPS server, which generates a high-precision 10 MHz reference clock through its internal clocking mechanism to synchronize the local clock of the camp control module. The precise GPS time, encoded in the IRIG-B code and synchronized with the one pulse per second (PPS) signal, is transmitted to the camp control module and then downwards the control module in the underground laboratory (hereafter referred to as lab control module) through a multi-kilometer optical fiber network. A hardware clock counter of the lab control module synchronizes with the DAQ’s master clock at the start of each new data acquisition run (labeled as local clock and local run start in Fig. 1 ), and the mapping between the delivered GPS time and the hardware clock ticks is stored in a database of a time server. The time delay introduced by the fiber optic length is calculated by measuring the time difference between the emission and reception of synchronized signals sent in opposite directions (downward and upward time streams), using the camp and lab control modules. This delay is utilized to correct timestamps during data acquisition. The precision of this system is maintained at the millisecond level, which is determined primarily by the GPS server.
Fig. 1: Schematic diagram of the PandaX-4T’s GPS-based timing system, which comprises a GPS receiver, a camp control module, a lab control module and a time server. The mapping of the delivered GPS time and the DAQ hardware clock ticks is stored in a database. See text for more details. Fig. 2: Schematic diagram of SN online monitoring system in PandaX-4T. The time consumed in each step is shown.
We developed an online software-based monitoring system to identify CCSN bursts by detecting transient increases in CENS-like event rate in PandaX-4T induced by SN neutrinos. The flow diagram is shown in Fig. 2. When the data acquisition system generates a new raw data file, the acquired data is stored using the Bamboo-shoot3 library [31, 32]. Each raw file contains approximately one minute of data, with a size of about 2 GB. The data compression process reduces it to about 1 GB and writes the data to local disk, which takes about ten seconds.
The data is then transmitted to the remote data center in Chengdu, enabling parallel processing of different files. The transmission time is approximately fifteen seconds. Offline processing and event reconstruction take about 4.5 minutes. After reconstruction, event information such as S1 and S2 variables, reconstructed vertex, and other relevant parameters is extracted. A series of cuts are then applied to reduce background, including the afterglow cut (which removes periods immediately following large signals amounting to a 11% of deadtime), the fiducial volume (FV) cut, the accidental coincidence (AC) cut, and quality cuts for S1 and S2.
[…]
Fig. 3: The background event rate (a) and the false alert (b) as a function of time.
Finally, a CCSN burst alert is generated if the number of neutrino candidates exceeds the threshold () within a 10-second time window. Further details on this method are provided in Ref. [35]. The SN alert information, including the alert start time and the total number of neutrino candidates, is sent to the PandaX-4T SN group via email. This process takes approximately five seconds. Considering all latency factors discussed above, the entire monitoring process takes about 6 minutes, as also indicated in Fig. 2.
The operational status of the SN monitoring system is displayed on a webpage, allowing personnel to inspect key performance parameters. The system operates stably during routine data-taking periods, with exceptions occurring only during special circumstances, such as power outages or server maintenance. SN monitoring program also automatically excludes calibration data based on the data type defined by the data acquisition system.
To evaluate the efficiency of the single event selection described above, we calculate the expected neutrino energy spectrum using the so-called Garching model with progenitor masses of and ( is the solar mass), employing the LS220 nuclear equation of state (EoS) [36, 37]. The electroweak neutral current process allows most neutrinos to interact coherently with nucleons in the nucleus for neutrino energies MeV. Details of the differential cross-section calculation can be found in Ref. [35].
@@
The signal response model for PandaX-4T is constructed using the Noble Element Simulation Technique (NEST) v2.3.6 parameterization [38], with relevant parameters tuned by fitting calibration data [39]. The NEST model simulates the light yield (, number of photons per keV) and charge yield (, number of electrons per keV) for nuclear recoil events in LXe. The S1 and S2 signals are then generated using a waveform simulation (WS) package, which assembles real waveforms from the data with and as inputs [40]. These WS events are reconstructed and selected using the same procedure as the real experimental data, allowing the overall detection efficiency to be determined.
The overall efficiency includes contributions from signal reconstruction, data quality selection, and the ROI cut, evaluated with the methodology outlined in Refs. [35, 41]. As an example, the overall detection efficiency for SN neutrino events using the Garching model with is approximately 13%. The most significant efficiency loss arises from the fact that the majority of CENS events have the low S1 and S2 values.
Based on the threshold mentioned above, eight alert candidates were observed during the 83 days of physics data. The time distribution of these alerts is shown in Fig. 3. Two alarms were identified in a same day, with a time difference of about 3 hours. No other anomalies were identified after a thorough check.
[…]
Based on this analysis, we conclude that within our 83-day dataset, the alerts are consistent with the absence of any real CCSN bursts above the background at the 1 level. It is important to note that the primary goal of SNEWS is to search for coincidences between multiple experiments, thereby significantly enhancing the sensitivity to CCSN signals beyond what is achievable by a single experiment. Notably, the false alert rate requirement for an experiment to join SNEWS is one per weeka criterion that PandaX-4T already satisfies.
[…]
Data type
Rate [/s]
Calendar time
Observed
Expected
DD
2.2
2.8 days
886
870.6
AmBe
9.7
3.7 days
288
265.3
Physical
2.9
83 days
8
6.1
For all eight alerts, intended alert messages to SNEWS were also generated, with one example alert shown in Table 2. More details about the SNEWS alert system can be found in Ref. [42].
For different (in total five) tiers of alerts, identified by the first column of Table 2, SNEWS provides different message template for participating experiments. For each tier, the second and third columns provide essential information for that example PandaX-4T alert.
Table 2: List of different message tiers within the SNEWS network, using a PandaX-4T alert as an example. The detector name is required in each tier and is not listed in the column.
For the Coincidence tier, the neutrino event time is represented in UTC format, derived from GPS time in PandaX-4T.
For the Significance tier, under the background only assumption, the p-values are the probabilities of observing more than or equal to the number of candidate events
within a 5-second time window at the single rate of .
For the Timing tier, the time series is an ordered sequence of nanosecond-precision integers, representing time offsets relative to the first neutrino candidate. For the Heartbeat tier, the detector status parameter is set to ON/OFF according to its operational condition.
For the Retraction tier, the retraction is initiated via a dedicated command provided by the SNEWS network, depending on real-time experimental considerations.
@@
Based on the performance of the SN monitoring system described above, the detection potential of various CCSN explosions with PandaX-4T-only data can be evaluated. The differential event rate of CENS can be expressed as
where represents the time after the CCSN core bounce, tonnes is the effective target mass in this analysis, is Avogadro’s number, and the molar mass of xenon is , and is the distance from the CCSN to Earth. represents the minimum neutrino incident energy required to induce a nuclear recoil energy of in CENS. In practice, we integrate over MeV). and are the differential cross-section of the CENS process and the total neutrino fluxes of all flavors at the energy of and time of , respectively. represents the detection efficiency.
The expected number of SN neutrino events in the detector can be calculated by integrating Eq. 1. Here, we integrate the variable for the model with ( and ) over the time interval from the onset time of core bounce to 8.35 (7.60) seconds. is integrated up to 30 keV. The results are summarized in Table 3. The two used typical distances of 10 kpc and 168 pc correspond to the positions of the nearby Galactic center and Betelgeuse [43] as observed from Earth, respectively. The detection probability of CCSN explosions, defined as the fraction of the observed over the total, is estimated using the Garching model with , employing the LS220 EoS. Figure 4 illustrates the detection probability as a function of CCSN distance to the PandaX-4T and the next generation experiment PandaX-20T, where we assume that PandaX-20T has a fiducial mass of tonnes. PandaX-4T achieves nearly- detection efficiency at 10 kpc, while the PandaX-20T maintains 100% efficiency up to 30 kpc.
Table 3: The number of expected SN neutrinos for source distances of 10 kpc and 168 pc in PandaX-4T, using Garching models ( and ) with the LS220 EoS. The fourth column indicates the time interval from the onset time of core bounce to the model’s termination time.
SN model
d=10 kpc
d=168 pc
Duration [s]
11.2
3.7
1.3
7.60
27
8.1
2.9
8.35
Fig. 4: Detection probability of the CCSN explosions as a function of distance using the Garching model with for PandaX-4T (black) and PandaX-20T (red). The corresponding false alert rate is twice a month in PandaX-4T.
Fig. 5: Typical timing distributions of the detected SN neutrino events from the distance of 168 pc (a) and 10 kpc (b). Data points are produced by simulations and the red lines are the fit results. The fitted onset times relative to the true time in about 10,000 mock detections are shown in the bottom panels. The resulting onset time measurement exhibits a variation of ms (c) and s (d) relative to zero for 168 pc and 10 kpc, respectively, within a 68% confidence level (C.L.).
The core bounce onset time can also be determined when sufficient event statistics are available. A typical fitting result of the timings after the core bounce of the SN neutrino events from the distance of 168 pc is displayed in Fig. 5. The points represent the simulated SN neutrino events, and the histogram represents the fitting result using the equation , where denotes the time profile function of the CCSN events, while and are fit parameters.
To estimate how well we can determine the onset time, signified by the precision on the parameter , we repeat the fits for many copies of mock time profiles, with results shown in Figs. 5 and 5. The 68% C.L. intervals for the variation of the offset times relative to the true value are ms at 168 pc and s at 10 kpc due to smaller statistics. Note that the described fitting method may not represent the timing precision achievable for promptly sharing data from a real supernova, as the true burst profile may not match the input model. As an alternative, we also consider the core bounce onset time as the arrival time of the first neutrino candidate, based on sampling from the Garching model. The resulting 68% C.L. intervals for the onset time variation are approximately 1 ms at 168 pc and 0.2 s at 10 kpc. The uncertainty at 168 pc is dominated by GPS timing precision.
In summary, we have established a robust online supernova monitoring system with GPS time synchronization implemented in PandaX-4T, leveraging the novel detection channel of coherent elastic neutrino-nucleus scattering. The performance of this system during an 83-day data-taking period, from July 18, 2025, to October 14, 2025, is presented. During this period, eight alerts were produced, consistent with a false alert rate of 2.2 per month caused by background. This monitoring system has the capability to provide the astronomical community with early warnings of CCSN explosions with a latency of about 6 minutes, and achieves nearly detection efficiency for supernovae occurring within 10 kpc of Earth in PandaX-4T. The system is ready to join the SNEWS network.
Looking ahead, the implementation of this system in the next-generation PandaX-20T experiment is expected to significantly enhance sensitivity to core-collapse supernovae within our galaxy. By enabling precise measurements of the recoil energy spectrum and time profile features, it offers an exciting opportunity to significantly advance our understanding of the fundamental properties of neutrinos and provide critical insights for refining astrophysical models.
@@
Acknowledgment: This project is supported in part by grants from National Key R&D Program of China (Nos. 2023YFA1606200, 2023YFA1606203), National Science Foundation of China (Nos. 12090060, 12090063, U23B2070, 12175139), and by Office of Science and Technology, Shanghai Municipal Government (grant Nos. 21TQ1400218, 22JC1410100, 23JC1410200, ZJ2023-ZD-003). We thank for the support by the Discipline Construction Fund of Shandong University, and the Fundamental Research Funds for the Central Universities. We also thank the sponsorship from the Chinese Academy of Sciences Center for Excellence in Particle Physics (CCEPP), Thomas and Linda Lau Family Foundation, New Cornerstone Science Foundation, Tencent Foundation in China, and Yangyang Development Fund. Finally, we thank the CJPL administration and the Yalong River Hydropower Development Company Ltd. for indispensable logistical support and other help.